Consider radium, which has a half-line of 1620 years. How many atoms of radium will decay in 5 s in 15 g of its sample? Consider the atomic weight of radium to be 226 kg/mol.

1
Expert's answer
2022-05-31T10:05:23-0400
"\Delta N=N_0-N=N_0(1-2^{-t/t_{1/2}})\\
=\frac{m}{\mu}N_A(1-2^{-t/t_{1/2}})""\Delta N=\frac{15}{226}*6.02*10^{23}(1-2^{-5/(1620*3.15*10^7)})\\
=4.0*10^{22}"